Find the slope of the line y= 3/4x.
Simplify your answer and write it as a proper fraction, improper fraction, or integer.
Thank you

Answer by addingup(3677)

The slope-intercept form is y=mx+b, where m is the slope and b the y-intercept
y= 3/4x
m = 3/4
b = 0
So the slope of your line is 3/4